# Verbalized Algorithms

### v3

@@ Abstract
Instead of querying LLMs in a one-shot manner and hoping to get the right answer for a reasoning task, we propose a paradigm we call verbalized algorithms (VAs), which leverage classical algorithms with established theoretical understanding. VAs decompose a task into simple elementary operations on natural language strings that they should be able to answer reliably, and limit the scope of LLMs to only those simple tasks. For example, for sorting a series of natural language strings, verbalized sorting uses an LLM as a binary comparison oracle in a known and well-analyzed sorting algorithm (e.g., bitonic sorting network). We demonstrate the effectiveness of this approach on sorting and clustering tasks.
@@ 1 Introduction
In contrast, traditional disciplines of computer science often approach such a task using task decomposition. For example, a merge sort decomposes the entire sorting task into several sorting tasks over individual sublists and combines the results, where, at the limit, there is an elementary binary comparison operator that compares individual pairs of objects in the list. These traditional algorithms, given an accurate oracle, guarantee the correctness of their outputs and the average/worst-case runtime.
@@
Such an algorithmic and analytical approach is often missing in LLM-based reasoning literature due to opaqueness of the reasoning mechanism in LLMs. More precisely, it often fails to distinguish the various separate challenges of reasoning tasks, i.e., (1) the ability to accurately perform atomic operations on ambiguous representations, (2) the ability to accurately execute a (not necessarily efficient) algorithm using those operations, (3) the choice of the algorithm to run and its theoretical properties (runtime/memory complexity, parallelism, robustness, soundness, completeness, etc.), and (4) the heuristic policy/hints for a branching algorithm. We argue that most reasoning tasks are informal representations of formal tasks for which an ideal classical algorithm with known complexity guarantees already exists in the literature, thus the effort to further improve (2,3) via additional training is wasteful and unproductive. For example, for the comparison sort, $O(\log n)$ algorithms are asymptotically optimal.
@@
In this work, we propose an abstract framework called “verbalized algorithms”. Given an informal representation of a formal task, a verbalized algorithm operates by replacing some basic atomic operations in a classical algorithm with LLM-driven text-based ones while retaining the high-level control flow of the classical algorithm, thereby largely retaining the original theoretical guarantees.
@@
This has advantages over another common approach to similar tasks that we call the “formalization-based approach”. Leveraging the machine translation ability of LLMs, it converts an informal task representation into a formal, symbolic representation, solves it with a specialized combinatorial solver, and then converts the solution back to the informal, natural language representation for the user. Existing work based on this approach includes LLM-based generation of PDDL (McDermott, 2000) for classical planning (Xie, 2020; Guan et al., 2023; Li et al., 2024; Fine-Morris et al., 2024; Oswald et al., 2024), Prover9 input language for theorem proving (Olausson et al., 2023), or SMT solver input for optimization (Hao et al., 2024). The issue with this approach is threefold: First, it requires grammar-following capability for a complex target formal language. Second, while the downstream solver requires a semantically and logically accurate input, the vagueness and informality in the natural language input often cause LLMs to generate inaccurate and inconsistent translations. Third, the assumptions made in the target formal language limits the scope of the tasks that can be solved by this approach. For example, a LLM-based planning agent that converts the task into PDDL cannot handle an informal task that contains probabilistic, temporal, or partially-observable nature. Verbalized algorithms, on the other hand, can handle them robustly via the general and informal common sense reasoning ability for atomic operations.
@@ 2 Verbalized Algorithms
Auto-regressive generative model is a probability distribution $p({\mathbf{x}})=p({\textnormal{x}}_{1},\ldots,{\textnormal{x}}_{N})$ where the random variable $\mathbf{x}$} represents a vector of integers, typically associated with a word or a sub-word token in natural language. Denoting its subsequence from index $i$ to $j$ as ${\mathbf{x}}_{i:j}$, Transformer-based language models (LMs) Vaswani et al. (2017) decompose it into each conditional distribution modeled by a large set of parameters $\theta$: $p({\mathbf{x}})=p_{\theta}({\mathbf{x}}_{N}|{\mathbf{x}}_{1:N-1})p({\mathbf{x}}_{1:N-1})=\ldots=\prod_{i=j}^{N-1}p_{\theta}({\mathbf{x}}_{i}|{\mathbf{x}}_{1:i-1})p({\mathbf{x}}_{1:j})$, where the prefix ${\mathbf{x}}_{1:j}$ is typically called a prompt and each conditional distribution $p_{\theta}({\mathbf{x}}_{i}|{\mathbf{x}}_{1:i-1})$ performs a next-token prediction.
To exhibit a trait typically useful for user interaction (e.g. instruction following), LLMs go through a series of parameter-efficient fine-tuning (PEFT) steps after pretraining on a large corpus. For example, in Direct Preference Optimization (Rafailov et al., 2023, DPO), the input is a pair of model outputs in which human users have indicated their preference over another and the model is fine-tuned to follow that preference. For a prompt $x$ and an answer pair $y_{1}$ and $y_{2}$, denote by $y_{1}\succ y_{2}|x$ that $y_{1}$ is preferred over $y_{2}$. DPO learns a Bradley-Terry reward model (Bradley and Terry, 1952, BT-model) for this preference, i.e., $p(y_{1}\succ y_{2}|x)$. The Plackett-Luce reward model (Plackett, 1975; Luce, 1959, PL-model) extends BT to a multinominal case that represents a ranking over choices.
@@ 3 Verbalized Sorting
Models (Qwen3) duplicates missing 1.7B 1.39 9.63 32B-AWQ 0.49 2.75
Table 2: Average Kendall-Tau scores over 5 seeds. Best scores in bold. The score of 1, 0, and -1 indicates perfect rank correlations, no correlations, and perfect anti-correlations, respectively.
@@
Table 5: Clustering scores across methods and benchmarks. The top 2 results are highlighted in bold.
method model $D=$ $k=$ Medarxiv IT Defects xkcdcolors-h xkcdcolors-l Jaccard 8 0.556 0.767 0.601 0.472 ST Qwen3-Emb…d-8B 4096 0.827 0.936 0.718 0.531 triplet Qwen3-4B 8 100 0.831 0.831 0.823 0.704 triplet Qwen3-4B 8 200 0.858 0.895 0.823 0.691 triplet Qwen3-32B-AWQ 8 100 0.842 0.908 0.851 0.748 triplet Qwen3-32B-AWQ 8 200 0.858 0.939 0.850 0.742
Verbalized Sorting sorts a list of natural language strings $X$ using LLM’s question answering capability as an atomic comparison oracle. Each binary comparison for $x_{1},x_{2}\in X$ is replaced by a yes-no style LLM query such as “Is X better than Y? X: $x_{1}$ Y: $x_{2}$”, where the next output token is restricted to “yes” and “no”. The model output is mapped to $\top/\bot$ if the sampled output is “yes” / “no”, respectively. This task is important in the RAG applications and also called the LLM-based ranking approach. We evaluated the accuracy of several non-VA and VA approaches. Method, dataset, and prompt details are available in the appendix Sec. A. All results are averages of 5 random seeds.
Constraint Decoding Baseline encodes the list $X$ in a markdown format and asks the model to return a sorted list using structured-decoding (Willard and Louf, 2023) to constrain the output to be a list whose elements are guaranteed to be one of the input strings. Due to the lack of a computationally efficient way to constrain the list to (1) be free of duplicates and (2) represent the same set of strings as the input, we (1) remove the duplicates and (2) append missing elements to the end of the list. The approach corresponds to listwise ranking approach Ma et al. (2023) in the RAG literature.
Scoring-based approaches ask the model to assign a score between 1.0 and 5.0 (down to one decimal) to each sentence using regex-constrained decoding, then sort the sentences using the scores. It is a form of formalization-based approach (the list of scores represents the formal problem). We tested two variants: I.i.d Scoring scores $x_{i}\in X$ independently, and Autoregressive Scoring scores all strings in a single query, allowing the model to see the scores of other elements. The approach corresponds to pointwise ranking approach Sun et al. (2023); Sachan et al. (2022) in the RAG literature.
VA approaches correspond to pairwise ranking approaches Qin et al. (2024) in the RAG literature, but our contribution focuses on leveraging the theoretical properties of classical algorithms, which has been undermined in the literature. VA Powersort overrides Python’s __gt__ operator with an LLM call and applies the built-in sorted function, which uses Powersort (Munro and Wild, 2018) that runs $O(n\log n)$ comparisons worst-case, the best runtime complexity achievable by comparison sort. VA Bitonic sorting network uses the Bitonic sorting network (Batcher, 1968) from parallel sorting literature to improve the throughput. It performs $O(n(\log n)^{2})$ comparisons, but runs in $O((\log n)^{2})$ time due to parallelism. As an additional optimization, we use the optimal sorting network (Knuth, 1997; Bundala and Závodnỳ, 2014) when the input size is 16 during its recursion. Robust VAs (Powersort/Bitonic) make individual comparisons more robust by sampling the output $K=3$ times as a batch LLM call and performing majority voting. For an odd $K$, if the oracle is correct with probability $p$, the voting result is correct with at least probability $1-e^{-K(2p-1)^{2}/2}$ (Emamjomeh-Zadeh and Kempe, 2018) due to Hoeffding bound (Hoeffding, 1963). Finally, Probabilistic VAs (Powersort/Bitonic) instead rely on comparing the probabilities / logits over the yes/no answers to perform each comparison query. While the naive probabilistic VA is equivalent to the naive VA with greedy sampling, we additionally have Symmetrized VA to address two biases in LLM: Positional bias and psycophancy (the tendency to reply “yes”) . Let $p_{12}$ and $p_{21}$ be next-token probabilities for answering “yes” with the original prompt and a prompt whose $x_{1}$ and $x_{2}$ are swapped. Due to the bias, $p_{12}\not=1-p_{21}$. Symmetrized VA tests $\frac{p_{12}+(1-p_{21})}{2}>\frac{(1-p_{12})+p_{21}}{2}$ as a comparison oracle.
@@
Next, Table 2 shows the average Kendall-Tau scores. Constraint-decoding baseline performed the worst overall, and while the scoring approaches can work with the large 32B model, their performance dwindles when the model is smaller. Robust VA with 32B achieved the best results. The strength of VA is most pronounced with the small 1.7B model, where other approaches failed. Notably, VAs with a 1.7B model outperformed the baseline with a 32B model, clearly showing the advantage of using the sorting algorithms developed in the literature.
Finally, Table 3 compares the runtime of each approach. VA Bitonic Sorting Network improves over the VA Powersort, and both VAs can be faster than the constrained-decoding baseline, even with robustification (x3 parallel queries) or symmetrization (x2 parallel queries). This shows that VAs can directly benefit from the theoretical results in the parallel sorting literature. Note that the parallel speedup in bitonic sort does not reach the theoretical limit (x11.4; 171 comparators and 15 steps for 25 inputs) because the maximum batch size is limited by the GPU; multi-GPU experiments with data-parallelism would improve this.
The Effect of Prefix Caching
Each query in VA sort requires a single next-token prediction. In Transformer architecture, it requires $O(t^{2})$ computations for $t$ input tokens. However, State-of-the-Art inference engines such as vllm Kwon et al. (2023) and SGLang Zheng et al. (2024) cache the prefix kv-cache over multiple queries. This is particularly helpful in reducing the inference cost of VA-style queries majority of which share the prefix.
@@
Assume that an input consists of $N+M$ tokens (prefix and suffix), and the kv-cache of $N$ prefix tokens is already available. Each $N+t$-th new input token attends to N+(t1)N+(t-1) existing tokens, which makes $N+M+1$-th token require $\sum_{t=1}^{M}(N+(t-1))=O(NM+M^{2})$ runtime.
With this in mind, assuming that the fixed prompt already has a kv-cache, the runtime of each comparison query reduces to $O(2L_{p}L+4L^{2})$, totalling $O(L_{p}^{2}+n^{2}(2L_{p}L+4L^{2}))$ runtime. Moreover, since $x_{1}$ is repeated $n$ times over $n^{2}$ comparisons, it further reduces to $O(L_{p}^{2}+n(L_{p}L+L^{2})+n^{2}((L_{p}+L)L+L^{2}))=O(L_{p}^{2}+n(L_{p}L+L^{2})+n^{2}(L_{p}L+2L^{2}))$. This is significant because the prompt tends to be much longer than each string.
Table 4 demonstrates the empirical speedup due to kv-cache in two State-of-the-Art inference engines. RadixAttention in SGLang is particularly helpful for VAs whose queries share large prefixes.

### v4

@@ Abstract
Instead of querying LLMs in a one-shot manner and hoping to get the right answer for a reasoning task, we propose a paradigm we call verbalized algorithms (VAs), which combines LLMs with classical algorithms with established theoretical guarantees. VAs decompose a task into simple elementary operations on natural language strings that LLMs are able to answer reliably, and limit the scope of LLMs to those simple tasks. For example, for sorting a series of natural language strings, verbalized sorting uses an LLM as a binary comparison oracle in a known and well-analyzed sorting algorithm (e.g., bitonic sorting network). Although this is already known as pairwise ranking in the literature, we additionally demonstrate the effectiveness of verbalized maximum, verbalized clustering, and verbalized submodular maximization for numerical reasoning, topic clustering and multi-hop Q&A RAG task, which guarantees $O(n)$ runtime, $O(n\log n)$ runtime, and 1/(1e)1/(1-e) optimality, respectively. Clustering and submodular maximization outperformed or improved the nearest neighbor search using state-of-the-art embedding models.
Large Language Model, Reasoning
[…]
@@ 1 Introduction
In contrast, traditional disciplines of computer science often approach such a task using task decomposition. For example, a merge sort decomposes the entire sorting task into smaller sorting tasks and combines the results, where, at the limit, there is an elementary binary comparison that compares an individual pair. These traditional algorithms, given an accurate oracle, guarantee the correctness of their outputs and the average/worst-case runtime/memory.
@@
Such an algorithmic and analytical approach is often missing in LLM-based reasoning literature due to opaqueness of the reasoning mechanism in LLMs. More precisely, it often fails to distinguish various separate challenges associated with reasoning, such as (1) the ability to accurately perform atomic operations on ambiguous representations, (2) the ability to accurately execute a sound but not necessarily efficient algorithm using those operations, (3) the theoretical efficiency of the algorithm (runtime/memory complexity, parallelism, robustness, soundness, completeness, etc.), and (4) the heuristic policy/hints for a branching algorithm. We argue that most reasoning tasks are informal representations of formal tasks for which an asymptotically optimal classical algorithm already exists in the literature, thus the effort to further improve (2,3) via additional training is wasteful and unproductive. For example, for comparison sort, $O(n\log n)$ algorithms are asymptotically optimal.
@@
In this work, we propose an abstract framework called “verbalized algorithms”. Given an informal representation of a formal task, a verbalized algorithm replaces some atomic operations in a classical algorithm with LLM-driven text-based oracles while retaining the high-level control flow, thereby largely retaining the theoretical guarantees.
@@
This has advantages over another common approach to similar tasks that we call the “formalization-based approach”. Leveraging the machine translation ability of LLMs, it converts an informal task representation into a formal, symbolic representation, solves it with a specialized combinatorial solver, and then converts the solution back to the informal, natural language representation for the user. Existing work based on this approach includes LLM-based generation of PDDL (McDermott, 2000) for classical planning (Xie, 2020; Guan et al., 2023; Li et al., 2024; Fine-Morris et al., 2024; Oswald et al., 2024), Prover9 input language for theorem proving (Olausson et al., 2023), or SMT solver input for optimization (Hao et al., 2024). The issue with this approach is threefold: First, it requires grammar-following capability for a complex target formal language. Second, while the downstream solver requires a semantically and logically accurate input, the vagueness and informality in the natural language input often cause LLMs to generate inaccurate and inconsistent translations. Third, the assumptions made in the target formal language limits the scope of the tasks that can be solved by this approach. For example, a LLM-based planning agent that converts the task into PDDL cannot handle an informal task that contains probabilistic, temporal, or partially-observable nature. Verbalized algorithms, on the other hand, can handle them robustly by applying a general and informal common sense reasoning to atomic operations.
This paper is organized as follows. We first formalize verbalize algorithms. We then perform four case studies with increasing difficulty and non-triviality: Maximum, sorting, clustering, and submodular maximization. Each section contains a task description, alternative algorithms considered, and an empirical evaluation. All experimental results are averages of 5 random seeds.
@@ 2 Verbalized Algorithms
Auto-regressive generative model is a probability distribution $p({\mathbf{x}})=p({\textnormal{x}}_{1},\ldots,{\textnormal{x}}_{N})$ where the random variable $\mathbf{x}$} represents a vector of integers, typically associated with a word or a sub-word token in natural language. Denoting its subsequence from index $i$ to $j$ as ${\mathbf{x}}_{i:j}$, Transformer-based language models (LMs) (Vaswani et al., 2017) decompose it into conditional distributions modeled by a large set of parameters $\theta$: $p({\mathbf{x}})=p_{\theta}({\mathbf{x}}_{N}|{\mathbf{x}}_{1:N-1})p({\mathbf{x}}_{1:N-1})=\ldots=\prod_{i=j}^{N-1}p_{\theta}({\mathbf{x}}_{i}|{\mathbf{x}}_{1:i-1})p({\mathbf{x}}_{1:j})$, where the prefix ${\mathbf{x}}_{1:j}$ is typically called a prompt and each conditional distribution $p_{\theta}({\mathbf{x}}_{i}|{\mathbf{x}}_{1:i-1})$ performs a next-token prediction.
To exhibit a trait typically useful for user interaction (e.g. instruction following), LLMs go through a series of parameter-efficient fine-tuning (PEFT) steps after pretraining on a large corpus. For example, in Direct Preference Optimization (Rafailov et al., 2023, DPO), the input is a pair of model outputs in which human users have indicated their preference over another and the model is fine-tuned to follow that preference. For a prompt $x$ and an answer pair $y_{1}$ and $y_{2}$, denote by $y_{1}\succ y_{2}|x$ that $y_{1}$ is preferred over $y_{2}$. DPO learns a Bradley-Terry reward model (Bradley & Terry, 1952, BT-model) for this preference, i.e., $p(y_{1}\succ y_{2}|x)$. The Plackett-Luce reward model (Plackett, 1975; Luce, 1959, PL-model) extends BT to a multinominal case that represents a ranking over choices.
@@
3 Verbalized Maximum
As a simplest algorithmic task to evaluate the reasoning accuracy of LLMs, we first investigate its ability to find the maximum number in a list $X$ of $n$ randomly selected integers between 1 and 10000. Due to the lack of principled algorithm for the LLM to follow, a naive LLM often fails even at this simple task despite with the help of constrained decoding, especially when the list gets longer.
We compare two approaches: Constraint Decoding Baseline queries LLMs with a string “What is the largest number in [comma-separated numbers in XX]?”. We use xgrammar constraint decoding engine to restrict the space of possible outputs to one of the elements in $X$.
Verbalized maximum overrides Python’s __gt__ operator with an LLM call and applies the built-in max function, which sequentially reduces the list with $O(n)$ comparisons. Each binary comparison for $x_{1},x_{2}\in X$ is replaced by a yes-no style LLM query “Is X larger than Y? X: $x_{1}$ Y: $x_{2}$”, where the next output token is restricted to “yes” and “no” with constrained decoding. The model output is mapped to $\top/\bot$ if the sampled output is “yes” / “no”, respectively.
We evaluated the algorithms above across 20 randomly sampled lists for each of $n\in{\left\{10,100\right\}}$ with various Qwen3 models (1.7B, 4B, 32B) with and without thinking enabled. We measured the mean absolute error (MAE) between the ground-truth maximum and the returned maximum. Note that the returned answer is guaranteed to be an element of $X$ in both algorithms.
The result is shown in Table 1. MAE column suggests that the baseline with a small 1.7B model struggles even on a small input $n=10$ and with the help of thinking. On larger inputs $(n=100)$, the accuracy degrades further on various model sizes, while verbalized maximum maintains a small, often zero error.
Table 1: Mean absolute error between the ground truth maximum and the LLM’s answer, and the runtime.
Models MAE($n=10$) MAE($n=100$) (Qwen3) baseline verbalized baseline verbalized 1.7B 500.14 0 1418.9 9.45 1.7B+think 14.18 0 1071.6 38.95 4B 0 0 34.25 0 4B+think 0 0 49.85 0 32B 0 0 48.6 0 32B+think 0 0 11.45 0
@@ 4 Verbalized Sorting
Verbalized Sorting sorts a list of natural language strings $X$ using LLM’s question answering capability as an atomic comparison oracle. Similar to the verbalized maximum, each binary comparison for $x_{1},x_{2}\in X$ is replaced by a yes-no style LLM query, although the exact prompt varies depending on the target application (see appendix Sec. A for details). This task is important in the RAG applications and also called the LLM-based ranking approach. We evaluated the accuracy of several non-VA and VA approaches.
Constraint Decoding Baseline encodes the list $X$ in a markdown format and asks the model to return a sorted list using structured-decoding (Willard & Louf, 2023) to constrain the output to be a list whose elements are guaranteed to be one of the input strings. Due to the lack of a computationally efficient way to constrain the list to (1) be free of duplicates and (2) represent the same set of strings as the input, we (1) remove the duplicates and (2) append missing elements to the end of the list. The approach corresponds to listwise ranking approach (Ma et al., 2023) in the RAG literature.
Scoring-based approaches ask the model to assign a score between 1.0 and 5.0 (down to one decimal) to each sentence using regex-constrained decoding, then sort the sentences using the scores. It is a form of formalization-based approach (the list of scores represents the formal problem). We tested two variants: I.i.d Scoring scores $x_{i}\in X$ independently, and Autoregressive Scoring scores all strings in a single query, allowing the model to see the scores of other elements. The approach corresponds to pointwise ranking approach (Sun et al., 2023; Sachan et al., 2022) in the RAG literature.
VA approaches correspond to pairwise ranking approaches (Qin et al., 2024) in the RAG literature, but our contribution focuses on leveraging the theoretical properties of classical algorithms, which has been undermined in the literature. VA Powersort overrides Python’s __gt__ operator with an LLM call and applies the built-in sorted function, which uses Powersort (Munro & Wild, 2018) that runs $O(n\log n)$ comparisons worst-case, the best runtime complexity achievable by comparison sort. VA Bitonic sorting network uses the Bitonic sorting network (Batcher, 1968) from parallel sorting literature to improve the throughput. It performs $O(n(\log n)^{2})$ comparisons, but runs in $O((\log n)^{2})$ time due to parallelism. As an additional optimization, we use the optimal sorting network (Knuth, 1997; Bundala & Závodnỳ, 2014) when the input size is 16 during its recursion. Robust VAs (Powersort/Bitonic) make individual comparisons more robust by sampling the output $K=3$ times as a batch LLM call and performing majority voting. For an odd $K$, if the oracle is correct with probability $p$, the voting result is correct with at least probability $1-e^{-K(2p-1)^{2}/2}$ (Emamjomeh-Zadeh & Kempe, 2018) due to Hoeffding bound (Hoeffding, 1963). Finally, Probabilistic VAs (Powersort/Bitonic) instead rely on comparing the probabilities / logits over the yes/no answers to perform each comparison query. While the naive probabilistic VA is equivalent to the naive VA with greedy sampling, we additionally have Symmetrized VA to address two biases in LLM: Positional bias (Liu et al., 2024) and psycophancy (the tendency to reply “yes”) (Casper et al., ). Let $p_{12}$ and $p_{21}$ be next-token probabilities for answering “yes” with the original prompt and a prompt whose $x_{1}$ and $x_{2}$ are swapped. Due to the bias, $p_{12}\not=1-p_{21}$. Symmetrized VA tests $\frac{p_{12}+(1-p_{21})}{2}>\frac{(1-p_{12})+p_{21}}{2}$ as a comparison oracle.
@@
Models duplicates missing Qwen3-1.7B 1.39 9.63 Qwen3-32B-AWQ 0.49 2.75
Table 3: Average Kendall-Tau scores over 5 seeds. Larger the better, best scores in bold.
@@
Next, Table 3 shows the average Kendall-Tau scores. Constraint-decoding baseline performed the worst overall, and while the scoring approaches can work with the large 32B model, their performance dwindles when the model is smaller. Robust VA with 32B achieved the best results. The strength of VA is most pronounced with the small 1.7B model, where other approaches failed. Notably, VAs with a 1.7B model outperformed the baseline with a 32B model, clearly showing the advantage of using the sorting algorithms developed in the classical discrete algorithm literature.
Finally, Table 4 compares the runtime of each approach. VA Bitonic Sorting Network improves over the VA Powersort, and both VAs can be faster than the constrained-decoding baseline, even with robustification (x3 parallel queries) or symmetrization (x2 parallel queries). This shows that VAs can directly benefit from the theoretical results in the parallel sorting literature. Note that the parallel speedup in bitonic sort does not reach the theoretical limit (x11.4; 171 comparators and 15 steps for 25 inputs) because the maximum batch size is limited by the GPU; multi-GPU experiments with data-parallelism would improve this result even further.
Effect of Prefix Caching
Each query in VA sort requires a single next-token prediction for yes/no. In Transformer architecture, it requires $O(t^{2})$ computations for $t$ input tokens. However, State-of-the-Art inference engines such as vllm (Kwon et al., 2023) and SGLang (Zheng et al., 2024) cache the prefix kv-cache over multiple queries. This is particularly helpful in reducing the inference cost of VA-style queries majority of which share the prefix.
@@
We analyze the runtime reduction due to kv-caches. Assume that an input query consists of $N+M$ tokens (prefix and suffix), and the kv-cache of $N$ prefix tokens is already available. Each $N+t$-th new input token attends to N+(t1)N+(t-1) existing tokens, which makes $N+M+1$-th token require $\sum_{t=1}^{M}(N+(t-1))=O(NM+M^{2})$ runtime. In other words, the runtime of each comparison query reduces to $O(2L_{p}L+4L^{2})$, totalling $O(L_{p}^{2}+n^{2}(2L_{p}L+4L^{2}))$ runtime. Moreover, since $x_{1}$ is repeated $n$ times over $n^{2}$ comparisons, it further reduces to $O(L_{p}^{2}+n(L_{p}L+L^{2})+n^{2}((L_{p}+L)L+L^{2}))=O(L_{p}^{2}+n(L_{p}L+L^{2})+n^{2}(L_{p}L+2L^{2}))$.
Although we assumed a simplified, exhaustive ($O(n^{2})$) comparison scenario, this speedup is significant in practice. Table 5 demonstrates the empirical speedup due to kv-cache in two state-of-the-art inference engines. RadixAttention in SGLang is particularly helpful for VAs whose queries share large prefixes.